Find value of: Integration of 1/x^2 dx with upper limit b and lower limit a.

Grade:11

1 Answers

AskiitianExpert Shine
14 years ago

Hi

1/(x^2) can be written as x^(-2)

And we know that the formula of integration for x^n (where n is not equal to -1) wrt x = (x^ (n+1)) / (n+1)

In our case, n = -2
Hence, the integration will give us (x^ (-1)) / (-1)
                                                      = -1/x
the upper and lower limits are given as b and a

This implies, integration = -1/b - (-1/a)
                                       = 1/a - 1/b